OpenStudy (anonymous):

You need $25,000 in 5 years. You invest 5% compound quarterly...how much do you need to put in to begin with?

OpenStudy (anonymous):

~$19,500

OpenStudy (anonymous):

Dude, he said compounding quarterly. 25000=x(1+.0125)^(5*4)

OpenStudy (anonymous):

Five years is twenty quarters. If interest is compounded quarterly, then the simple interest rate is 1.25% per quarter. Let x be the initial deposit. If nothing is withdrawn of deposited during the next five years, then the amount of the account at the end of feve years is \[x(1.025)^{20} = $25,000.\]Solve for x to get \[x = $25,000/(1.025)^{20} = $15,256.77\]

OpenStudy (anonymous):

^You have an error in your calculations - you used 1.025 instead of 1.0125. That's why your answer is ~$4000 off. o.O

OpenStudy (anonymous):

You're right about that. My glitch

OpenStudy (anonymous):

so \[$25,000/(1.0125)^{20} = $19,500.21.\]

OpenStudy (anonymous):

That's much better.